The molar volume of isobutane at 300.0 K and 1 bar is 24.31 dm3/mol. Estimate the value of B2V for isobutane at 300K.
Answer #1

Given that, temperature (T) = 300 K

Pressure (P) = 1 Bar

Gas constant (R) = 0.083145 dm^3 bar /mol K

Let V is the molar volume in ideal condition. Then using ideal gas equation for 1 mole,

PV = RT

V = RT/P = 0.083145 * 300 / 1 = 24.944 dm^3

B2V value = Given molar volume - ideal molar volume = 24.31 - 24.944 = -0.634 dm^3/mol .... Answer